# Network Formation and Dynamics Among Multi-LLMs

### v2

@@ 2.2. Macro-Level Principles
Next, we use modularity maximization (Blondel et al., 2008) to examine the community structure that arises due to homophily. Intuitively, the modularity measure compares the number of edges inside a community with the expected number of edges that one would find in the cluster if the network were a random network with the same number of nodes and where each node keeps its weighted degree, but edges are otherwise randomly attached according to the Chung-Lu model (Chung and Lu, 2004). This random null model implicitly assumes that each node can attach to another network node. Formally, for a graph with weights $w_{ij}$ on its edges, the weighted modularity $Q$ (Clauset et al., 2004) of a partition with $C$ communities given by
$Q=\sum_{c=1}^{C}\left[\frac{L_{c}}{W}-r\left(\frac{k_{c}}{2W}\right)^{2}\right],$
Here $W$ is the sum of all weights, $L_{c}$ is the total weight of intra-community links for community $c$, $k_{c}$ is the total weighted degree of the nodes in the community $c$ and $r$ is the resolution parameter which (we take $r=1$ from now on). A large positive value (such as greater than 0.5) of the modularity corresponds to evidence of community structure, corresponding to deviation from the random null model. Maximizing $Q$ is NP-Hard, we use the Louvain algorithm (Blondel et al., 2008) to get a lower bound on the maximum modularity.
In Figure 3(b), we illustrate the communities identified by the Louvain algorithm at various temperatures. Alongside, we present the average Louvain modularity values over ten simulations, with 95% confidence intervals (refer to Figure 3(c)). For the network’s weights, we use the number of common attributes shared between each pair of nodes. This is defined as $w_{ij}=\left|\left\{k:x_{i}^{(k)}=x_{j}^{(k)}\right\}\right|$ for each link $(i,j)$ in the final network. Here, $x_{i}^{(k)}$ and $x_{j}^{(k)}$ correspond to the $k$-th features of $x_{i}$ and $x_{j}$, respectively.
We observe that the Louvain partitions demonstrate community structure with positive modularity across all temperatures. This is confirmed by a t-test comparing the modularity to $Q=0$ of a random graph ($P<0.001$). The modularity is most pronounced at lower temperatures. As the temperature increases, the Louvain modularity also increases ($P<0.001$, t-test). Furthermore, lower temperatures result in fewer (and larger) communities on average, as identified by the Louvain algorithm ($P<0.001$, t-test). Additionally, communities tend to be visually more distinct at lower temperatures. We propose that higher temperatures reflect increased randomness in the decisions.
Our results demonstrate that community structures manifest in networks generated by LLMs. Specifically, these networks adhere to the principle of triadic closure, which further reinforces the formation of community structures. Additionally, LLM-generated networks exhibit homophily, facilitating the emergence of distinct communities. In Section 2.3, we further illustrate that community structures are reinforced in real-world networks, where LLMs consider a combination of factors, including homophily and triadic closure, in their decisions to form connections.
@@
(d) Estimated $\hat{\beta}$ for $\beta=0.25$
(e) Estimated $\hat{\beta}$ for $\beta=0.5$
(f) Estimated $\hat{\beta}$ for $\beta=0.75$
Figure 4. Figures 4(a), 4(b) and 4(c): and distribution of reasoning created based on Principle 5 using the altered Watts-Strogatz Model for $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$, together with plots of the average clustering coefficient $C$ and the average shortest path length $L$. The comparison is made with respect to a Watts-Strogatz graph with $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$. In most cases, we observe strong evidence that $L\sim\log(n)$ and $C\sim 1/\log(n)$. Figures 4(d), 4(e) and 4(f): Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with rewiring probability $\hat{\beta}$. We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$.
@@
Formally, a small-world network is a network where the average shortest path length $L$ grows proportionally to the network size $n$, [^8] This follows the definition in (Jacobs and Watts, 2021). namely
@@
We employ a modified version of the model, where edge rewiring is informed by LLM queries, based on the current network structure. The generation process is parametrized by the number of nodes ($n$), the number of long-range connections ($k$), and the rewiring probability ($\beta$). The algorithm proceeds as follows:
@@
We generate networks with varying sizes ($n=10$ to $n=100$) and examine the relationships between the average shortest path length ($L$) and the average clustering coefficient ($C$) with respect to the network size ($n$). We test $\beta\in\{0.25,0.5,0.75\}$ and set $k=5$ as an illustration. In addition, we compare with the results reported in (Jacobs and Watts, 2021). Some instances of these networks are shown in Figures 4(a), 4(b) and 4(c). We run regressions relating to the average shortest path length with $\log(n)$ and observe that for all temperatures, we obtain statistically significant results (see the significance indicators in Figure 5, where most regresssions have $P<0.001$) being that the average shortest path length scales proportionally to $\log(n)$. Similarly, we show that the average clustering coefficient scales as $1/\log(n)$ (again, most regression results have $P<0.001$). Both results follow the guideline from (Jacobs and Watts, 2021) which tested the small-world properties of organizational networks.
(a) $\beta=0.25$
(b) $\beta=0.5$
(c) $\beta=0.75$
Figure 5. Regression plots relating average shortest path length and average clustering coefficient with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$. $a$ represents the slope of the regression lines. *** indicates $P<0.05$, ***** * indicates $P<0.01$, and ******* * * indicates $P<0.001$.
However, a direct comparison with the Watts-Strogatz model reveals that the LLM-generated networks do not exactly mirror the statistics of Watts-Strogatz networks with corresponding $\beta$ values. As Figures 4(a), 4(b) and 4(c) show, we only observe weak evidence that the final generated networks have a similar average shortest path length as Watts-Strogatz with $\beta\in\{0.25,0.5,0.75\}$ ($P<0.1$, t-test comparing the resulting average shortest path lengths of the networks generated by multi-LLMs and the Watts-Strogatz average shortest path lengths for $\beta\in\{0.25,0.5,0.75\}$) and a larger average clustering coefficient than Watts-Strogatz ($P<0.1$, t-test comparing the resulting average clustering coefficient of the networks generated by multi-LLMs and the Watts-Strogatz average clustering coefficient for $\beta\in\{0.25,0.5,0.75\}$).
To quantify this resemblance, we conduct a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks. The results, depicted in Figures 4(d), 4(e) and 4(f), show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings. In Figures 4(d), 4(e) and 4(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$, each simulation and each temperature together with the $P$-values resulting from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$.
In conclusion, our analysis demonstrates that LLM-generated networks exhibit key small-world properties, with logarithmic scaling of average shortest path lengths and inverse logarithmic scaling of average clustering coefficients. While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics, underscoring the potential of LLMs in generating small-world networks.
2.3. Real-World Networks
To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972). This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network. We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives.
2.3.1. Discrete Choice Model
@@
In this equation, $\theta_{\mathrm{PA}}$ measures the strength of preferential attachment based on the degree $d_{j,t}$ of $j$ at step $t$, $\theta_{\mathrm{H}}$ measures the strength of homophily based on the similarity $w_{ij}$ (i.e. number of common attributes) between $i$ and $j$, and $\theta_{\mathrm{TC}}$ measures the strength of triadic closure, based on the number of common neighbors $c_{ij,t}$ between $i$ and $j$ at step $t$. The error term $\epsilon_{ij,t}$ is distributed as i.i.d. standard Gumbel [^9] The standard Gumbel distribution has CDF $e^{e^{-x}}$.. All variables are first normalized based on their range, and then the log transformation is taken.
@@
2.3.2. Experiment Setup and Rationale
We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30. They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively.
For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the nodes of the networks as nodes that have to make a choice. For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models. We set the size of the alternative sets $A_{t}$ to 20, and thus, the accuracy of random guessing is 5%.
Figure 6. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown with error bars. The full regression tables can be found in Appendix A. All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
2.3.3. Results
We present the regression results in Figure 6101010More detailed results can be found in Table S.1 in Appendix A. Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets. Furthermore, we find that LLMs place the greatest emphasis on the number of common attributes ($\hat{\theta}_{\mathrm{H}}$), followed by the number of common neighbors ($\hat{\theta}_{\mathrm{TC}}$), and then the degree ($\hat{\theta}_{\mathrm{PA}}$).
In addition to the coefficient $\hat{\theta}_{\mathrm{TC}}$ of triadic closure, we measure the percentage change in the average shortest path length and the average clustering coefficient between the network before the edge deletions and the network after the choices made by the LLM agents (see Table S.1). We observe that the average clustering coefficient also decreases compared to the initial network, indicating the weakening of triadic closure compared to the initial network.
We also provide reasoning analysis. Figure S.8 demonstrates that the rationales provided by LLMs aligns with the aforementioned findings. In particular, LLMs emphasize homophily, indicating that nodes sharing similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to form connections. Conversely, the effects of triadic closure and preferential attachment are less prominent.
We also compare LLMs’ prediction with random guessing under a link prediction framework. We evaluate the percentage increase in accuracy, where accuracy is defined as the likelihood of selecting a previously existing neighbor from a set of alternatives, compared to making a random selection from 20 possible options. As shown in Appendix A, LLMs demonstrate a substantially higher accuracy in their selections than would be expected from random guessing.
Moreover, we examine whether the community structure is reinforced after the network formation decision made by LLMs. We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularities (see Appendix A). We observe that the modularity increases in all cases and is statistically significant. This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
To reflect the relation with “small-world” phenomenon, we measure the percent change in the average shortest path length between the network before the edge deletions and the network after the choices made by the LLM agents (see Appendix A). We observe a small decrease in the average path length in the new network in almost all cases indicating that the average shortest path length decreases slightly, indicating strengthening of the small-world phenomenon.
Taken together, our analysis indicates that although preferential attachment, homophily, and triadic closure are all present in the network formation behaviors of LLMs, homophily emerges as the predominant driving force. Moreover, both community structure and small-world phenomenon are reinforced.
3. Discussion
Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon. Our findings indicate that networks generated by multiple LLMs do exhibit these properties, particularly when primed with network statistics such as the number of mutual friends or the degree of alternatives. Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks. Our results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment. Additionally, we observed a strengthening of community structure in real-world network experiments.
Furthermore, our study may pave the way for innovative applications of LLMs in network science research. Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling. LLM agents have the capability to produce more sophisticated and nuanced behaviors, including negotiation, collaboration, and competition. This opens up opportunities to simulate networked environments such as marketplaces or organizational structures, where LLMs can model decision-making processes and interaction patterns.
4. Methods
4.1. Experimental Procedure
Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$. The initial network, $G_{1}$, is referred to as the seed network. At each step $t$, we select a query node $i_{t}$ (which may either be a new arrival or an existing node in the graph) and assign it the task of forming new links. This is accomplished by selecting nodes from a set of alternatives $A_{t}$ (meaning potential candidates for link formation) and initiating a query call $\mathcal{Q}(A_{t},i_{t},\delta)$ to the LLM (as outlined in Algorithm 1) to create up to $\delta$ new links. The edge set selection process involves presenting the LLMs with personal or network features of the alternatives, denoted as $F(A_{t})=\left\{f_{a}:a\in A_{t}\right\}$, which may include information such as the neighbors of the nodes, node degrees, common connections with $i_{t}$, and community memberships, formatted in JSON. We adopt a zero-shot learning approach, avoiding the provision of examples to the model to prevent bias, in line with relevant studies such as (Brookins and DeBacker, 2023). This approach allows for the exploration of the innate preferences of LLMs.
4.2. Summarizing the Reasoning of the LLM Agents
In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions. Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices. We then aggregate all the provided rationales using GPT models (GPT-3.5 for Section 2.1 and GPT-4 for Section 2.3) by asking it to identify the common themes, resulting in the creation of multiple categories. To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
4.3. Estimating the Parameters of the Discrete Choice Model
@@
$(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})=\operatorname*{arg\,max}_{(\theta_{\mathrm{PA}},\theta_{\mathrm{TC}},\theta_{\mathrm{H}})\in\mathbf{R}^{3}}\sum_{t=1}^{T}\left(\theta_{\mathrm{PA}}\log d_{j_{t},t}+\theta_{\mathrm{H}}\log w_{i_{t}j_{t}}+\theta_{\mathrm{TC}}\log c_{i_{t}j_{t},t}-\log\left(\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{i_{t}r}^{\theta_{\mathrm{H}}}c_{i_{t}r,t}^{\theta_{\mathrm{TC}}}\right)\right),$

### v3

@@ Results
Next, we investigate the community structure resulting from homophily using modularity maximization (Blondel et al., 2008) (Figure 3). Modularity quantifies the discrepancy between the actual number of edges within communities and the expected number in a random network with identical node count and degree distribution, following the Chung-Lu model (Chung and Lu, 2004). This model presumes that nodes maintain their weighted degree, with edges randomly distributed. The weighted modularity $Q$ (Clauset et al., 2004) for a graph with edge weights $w_{ij}$ and $C$ communities is defined as
$Q=\sum_{c=1}^{C}\left[\frac{L_{c}}{W}-r\left(\frac{k_{c}}{2W}\right)^{2}\right].$
Here $W$ represents the total edge weights, $L_{c}$ the intra-community link weights for community $c$, $k_{c}$ the total weighted degree within community $c$, and $r$ the resolution parameter, set to 1 for our analysis. High modularity values (e.g., greater than 0.5) indicate significant community structuring, diverging from the random model. [^8] Given the NP-Hard nature of maximizing $Q$, we employ the Louvain algorithm (Blondel et al., 2008) to approximate the highest possible modularity.
For the network’s weights, we use the number of common attributes shared between each pair of nodes: $w_{ij}=\left|\left\{k:x_{i}^{(k)}=x_{j}^{(k)}\right\}\right|$ for each link $(i,j)$ in the final network. Here, $x_{i}^{(k)}$ and $x_{j}^{(k)}$ correspond to the $k$-th features of $x_{i}$ and $x_{j}$, respectively.
In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures. Notably, communities appear more distinct at lower temperatures, likely due to reduced randomness in decision-making at these temperatures. Additionally, Figure 3(c) presents the distribution of Louvain modularity values across simulations, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$). Moreover, lower temperatures lead to fewer, larger communities as per the Louvain algorithm’s findings ($P<0.001$, t-test).
Our results demonstrate that community structures manifest in networks generated by LLMs, driven by both triadic closure and homophily.
@@
Figure 4. Simulation results for Principle 5 (small world). Network instances for the networks created according to Principle 5 using the altered Watts-Strogatz Model for node count $n=50$, average degree $k=5$, rewriting probability $\beta\in\{0.25,0.5,0.75\}$, together with plots of the average clustering coefficient $C$ and the average shortest path length $L$. The comparison is made with respect to a Watts-Strogatz graph with $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$. The error bars correspond to 95% confidence intervals.
(a) $\beta=0.25$
(b) $\beta=0.5$
(c) $\beta=0.75$
(d) $\beta=0.25$
(e) $\beta=0.5$
(f) $\beta=0.75$
Figure 5. Fitted results for Principle 5 (small world). Top: Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$. Value $a$ in legends represents the slope of the regression lines. Bottom: Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$. We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$. *: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.
@@
A small-world network is defined by its average shortest path length $L$, which grows logarithmically with the size of the network $n$, [^9] As per the definition in (Jacobs and Watts, 2021). expressed as
@@
We employ a modified version of the model, where edge rewiring is informed by LLM queries, based on the current network structure. The generation process is parametrized by the number of nodes ($n$), average degree ($k$), and the rewiring probability ($\beta$). See details in Methods and Materials.
We generated networks of various sizes, ranging from $n=10$ to $n=100$, to explore the relationship between the network size ($n$) and two key metrics: the average shortest path length ($L$) and the average clustering coefficient ($C$). For this analysis, we considered values of $\beta$ set at 0.25, 0.5, and 0.75, with a fixed $k=5$ to serve as a consistent parameter. Visual representations and average clustering coefficients are presented in Figures 4(a), 4(b) and 4(c).
However, when directly compared with the Watts-Strogatz model, the networks generated by the LLMs do not precisely replicate the characteristics of Watts-Strogatz networks for the corresponding the rewiring probabilities ($\beta$). As illustrated in right panels of Figures 4(a), 4(b) and 4(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths. Additionally, LLM-generated networks exhibit a larger average clustering coefficient than those of the Watts-Strogatz model for the same rewiring probabilities $\beta$, also with a significance level of $P<0.1$ in the t-test comparisons. This discrepancy suggests the differences in the network structure and connectivity patterns between the LLM-generated networks and the classical Watts-Strogatz model.
We also provide regressions analysis by examining the correlation between the average shortest path length and average clustering coefficient versus $\log(n)$ (refer to Figure 5(a), Figure 5(b), and Figure 5(c)). We found that across all tested temperatures, the relationships were statistically significant, with most regressions yielding $P<0.001$. This indicates that the average shortest path length increases proportionally with $\log(n)$. Similarly, for the average clustering coefficient, we demonstrated that it inversely scales with $1/\log(n)$, with the majority of regression analyses also showing $P<0.001$. These findings align with the small-world properties of organizational networks as documented in the study by (Jacobs and Watts, 2021), suggesting that these characteristics are not only prevalent but also predictable across different network sizes.
To quantify how LLM-generated networks resemble Watts-Strogatz networks, we fit the estimated $\hat{\beta}$ values for each LLM-generated network. [^10] We conducted a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks. In Figures 5(d), 5(e) and 5(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature. Here $P$-values result from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$. These results show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings.
In conclusion, our analysis demonstrates that LLM-generated networks exhibit key small-world properties, with logarithmic scaling of average shortest path lengths and inverse logarithmic scaling of average clustering coefficients. While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics.
Decisions on Real-World Networks
To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972). This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network. We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30. They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively. For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the network nodes to be nodes that have to make a choice. For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models. We set the size of the alternative sets $A_{t}$ to 20, and thus, the accuracy of random guessing is 5%.
Figure 6. Regression results for real-world network experiments. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown with error bars. The full regression tables can be found in SI Section 1. All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
We regress network formation decisions on standardized scores reflecting the three micro-level principles. We present the regression results in Figure 6.111111More detailed results can be found in Table S1 in SI Section 1. Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets. Furthermore, we find that LLMs place the greatest emphasis on the number of common attributes ($\hat{\theta}_{\mathrm{H}}$), followed by the number of common neighbors ($\hat{\theta}_{\mathrm{TC}}$), and then the degree ($\hat{\theta}_{\mathrm{PA}}$).
Further insights are provided in SI Section 4, where we provide the reasoning analysis. The analysis points to a marked emphasis on homophily, indicating that nodes with similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to connect. In contrast, triadic closure and preferential attachment have a relatively lesser impact.
Our study also includes a comparison of LLMs’ predictive capabilities against random guessing in a link prediction framework. We assess the percentage increase in accuracy, defining accuracy as the probability of selecting an existing neighbor from a set of alternatives, as opposed to a random selection from 20 possible options. As detailed in SI Section 1, LLMs significantly outperform random guessing in accuracy.
Furthermore, we investigate whether LLMs reinforce community structure post-network formation. We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularity scores (see SI Section 1). We observe that the modularity increases in all cases and is statistically significant. This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
Additionally, to explore the “small-world” phenomenon, we measure the percent change in average shortest path length between the pre- and post-decision networks (detailed in SI Section 1). Our findings reveal a minor decrease in average path length in most cases, suggesting a slight strengthening of the small-world phenomenon.
In summary, our analysis demonstrates that while homophily, triadic closure, and preferential attachment are integral to the network formation behaviors of LLMs, homophily is the dominant factor. Furthermore, both community structure and the small-world phenomenon are reinforced after network formation decisions of LLMs.
Discussion
Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon. Our findings indicate that networks generated by multiple LLMs do exhibit these properties, particularly when primed with network statistics such as the number of mutual friends or the degree of alternatives. Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks. Our real-world network results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment. Additionally, we observed a strengthening of community structure in real-world network experiments.
Furthermore, our study may pave the way for innovative applications of LLMs in network science and computational social science research. Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling. As LLM agents gain increasing ability to exhibit more complex and nuanced behaviors, such as negotiation, collaboration, and competition, this presents opportunities to replicate networked environments like marketplaces or organizational structures. In these settings, LLMs could effectively model decision-making processes and interaction dynamics.
1. Methods
Methods and Materials
ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
Experimental Procedure
Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$ with vertex sets $V_{1},\dots,V_{T}$. The initial network, $G_{1}$, is referred to as the seed network. At each step $t$, we select a query node $i_{t}$ (which may either be a new arrival or an existing node in the graph) and assign it the task of forming new links. This is accomplished by selecting nodes from a set of alternatives $A_{t}$ (meaning potential candidates for link formation) and initiating a query call $\mathcal{Q}(A_{t},i_{t},\delta)$ to the LLM (as outlined in Algorithm 1) to create up to $\delta$ new links. The edge set selection process involves presenting the LLMs with personal or network features of the alternatives, denoted as $F(A_{t})=\left\{f_{a}:a\in A_{t}\right\}$, which may include information such as the neighbors of the nodes, node degrees, common connections with $i_{t}$, and community memberships, formatted in JSON. We adopt a zero-shot learning approach, avoiding the provision of examples to the model to prevent bias, in line with relevant studies such as (Brookins and DeBacker, 2023). This approach allows for the exploration of the innate preferences of LLMs.
Summarizing the Reasoning of the LLM Agents
In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions. Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices. We then aggregate all the provided rationales using GPT models (GPT-3.5 for the synthetic networks and GPT-4 for the real-world networks) by asking it to identify the common themes, resulting in the creation of multiple categories. To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
Details for Small-World Experiments
The algorithm of generating LLM networks is described as follows:
@@ 1. Methods
The Discrete Choice Model in Real-World Network Experiments
@@
In this equation, $\theta_{\mathrm{PA}}$ measures the strength of preferential attachment based on the degree $d_{j,t}$ of $j$ at step $t$, $\theta_{\mathrm{H}}$ measures the strength of homophily based on the similarity $w_{ij}$ (i.e. number of common attributes) between $i$ and $j$, and $\theta_{\mathrm{TC}}$ measures the strength of triadic closure, based on the number of common neighbors $c_{ij,t}$ between $i$ and $j$ at step $t$. The error term $\epsilon_{ij,t}$ is distributed as i.i.d. standard Gumbel. [^12] The standard Gumbel distribution has CDF $e^{e^{-x}}$. All variables are first normalized based on their range, and then the log transformation is taken.
@@
Estimating the Parameters of the Discrete Choice Model
@@
$(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})=\operatorname*{arg\,max}_{(\theta_{\mathrm{PA}},\theta_{\mathrm{TC}},\theta_{\mathrm{H}})\in\mathbf{R}^{3}}\sum_{t=1}^{T}\bigg{(}\theta_{\mathrm{PA}}\log d_{j_{t},t}+\theta_{\mathrm{H}}\log w_{i_{t}j_{t}}+\theta_{\mathrm{TC}}\log c_{i_{t}j_{t},t}-\log\big{(}\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{i_{t}r}^{\theta_{\mathrm{H}}}c_{i_{t}r,t}^{\theta_{\mathrm{TC}}}\big{)}\bigg{)},$
@@
Data and Code Availability
Data and code are available at GitHub: https://github.com/papachristoumarios/llm-network-formation.